1. An object moving in a straight line at 15 m/s is under constant acceleration. After 5.0 s, its velocity is 17.5 m/s.

a.How far did the object travel this time?


b.What is the acceleration of the object?



1
Expert's answer
2021-10-31T12:42:45-0400

at t = 0.0s, v = u = 15m/s

at t = 5.0s, v = 17.5m/s


a.

"s = (\dfrac{u+v}2)t= (\dfrac{15+17.5}2)×5\\
s =16.25×5 = 81.25m"


b.

"v = u+at \\
17.5 =15+ a(5)\\
2.5 = 5a\\
a = 0.5m/s²"